Sammy Slick works for a company that allows him to contribute up to 10% of his earnings into a tax-deferred savings plan. The company matches a portion of the contributions its employees make based on the organization's financial performance. Although the minimum match is 25% of the employee's contributions and the maximum match is 100%, in most years the company match is about 50%. Sammy is currently 30 years old and makes $35,000. He wants to retire at age 60. He expects his salary to increase in any given year to be at least 2% per year, at most 6%, and most likely 3.5%. The funds contributed by Sammy and his employer are invested in mutual funds. Sammy expects the annual return on his investments to vary according to a normal distribution with a mean of 12.5% and standard deviation of 2%.

a. If Sammy contributes 10% of his income to this plan, how much money could he expect to have at age 60? Round your answer to three decimal places.
$  million

b. Suppose Sammy makes 10% contributions to this plan for eight years, from age 30 to 37, and then stops contributing. How much of his own money would he have invested and how much money could he expect to have at age 60? Round your answers to three decimal places, if necessary.

Total investment $ million
Ending balance $ million

c. Now suppose Sammy contributes nothing to the plan his first eight years and then contributes 10% for twenty-three years from age 38 to age 60. How much of his own money would he have invested and how much money could he expect to have at age 60? Round your answers to three decimal places, if necessary.

Total investment $ million
Ending balance $ million
Answer #1

ANSWER;

Number of years to retire 30 (60-30)
Current Salary $35,000
Next years expected salary $36,225 (35000*1.035)
Salary in year (N+1)=1.035*Salary in year( N)
First years saving $3,500 (35000*0.1)
Expected matching contribution by the company $1,750 (3500*0.5)
Mean return from investment=12.5% 0.125
Future Value(FV) of Savings=(Savings)*((1+i)^(30-N))
i=expected return from investment=0.125
N=Year of Saving
Year wise Saving and FV of savings are given below

A

N A B=A*0.1 C=0.5*B D=B+C E=D*((1.125^(30-N)
Year Expected Salary Expected Saving Expected Match Total Saving FV of Total Saving
1 $35,000 $3,500 $1,750 $5,250 159802.0898
2 $36,225 $3,623 $1,811 $5,434 147017.9226
3 $37,493 $3,749 $1,875 $5,624 135256.4888
4 $38,805 $3,881 $1,940 $5,821 124435.9697
5 $40,163 $4,016 $2,008 $6,024 114481.0921
6 $41,569 $4,157 $2,078 $6,235 105322.6047
7 $43,024 $4,302 $2,151 $6,454 96896.79636
8 $44,530 $4,453 $2,226 $6,679 89145.05265
9 $46,088 $4,609 $2,304 $6,913 82013.44844
10 $47,701 $4,770 $2,385 $7,155 75452.37256
11 $49,371 $4,937 $2,469 $7,406 69416.18276
12 $51,099 $5,110 $2,555 $7,665 63862.88814
13 $52,887 $5,289 $2,644 $7,933 58753.85708
14 $54,738 $5,474 $2,737 $8,211 54053.54852
15 $56,654 $5,665 $2,833 $8,498 49729.26464
16 $58,637 $5,864 $2,932 $8,796 45750.92347
17 $60,690 $6,069 $3,034 $9,103 42090.84959
18 $62,814 $6,281 $3,141 $9,422 38723.58162
19 $65,012 $6,501 $3,251 $9,752 35625.69509
20 $67,288 $6,729 $3,364 $10,093 32775.63948
21 $69,643 $6,964 $3,482 $10,446 30153.58833
22 $72,080 $7,208 $3,604 $10,812 27741.30126
23 $74,603 $7,460 $3,730 $11,190 25521.99716
24 $77,214 $7,721 $3,861 $11,582 23480.23739
25 $79,916 $7,992 $3,996 $11,987 21601.8184
26 $82,714 $8,271 $4,136 $12,407 19873.67292
27 $85,609 $8,561 $4,280 $12,841 18283.77909
28 $88,605 $8,860 $4,430 $13,291 16821.07676
29 $91,706 $9,171 $4,585 $13,756 15475.39062
30 $94,916 $9,492 $4,746 $14,237 14237.35937
SUM 1833796.489
Sammy is expected to have at the age of 60 $ 1,833,796.49

Sammy is expected to have at the age of 60

$1.834 million

B

If Sammy has contributed for 8 years from 30 to 37 years
N A B=A*0.1 C=0.5*B D=B+C E=D*((1.125^(30-N)
Year Expected Salary Expected Saving Expected Match Total Saving FV of Total Saving
1 $35,000 $3,500 $1,750 $5,250 159802.0898
2 $36,225 $3,623 $1,811 $5,434 147017.9226
3 $37,493 $3,749 $1,875 $5,624 135256.4888
4 $38,805 $3,881 $1,940 $5,821 124435.9697
5 $40,163 $4,016 $2,008 $6,024 114481.0921
6 $41,569 $4,157 $2,078 $6,235 105322.6047
7 $43,024 $4,302 $2,151 $6,454 96896.79636
8 $44,530 $4,453 $2,226 $6,679 89145.05265
SUM $31,681 SUM $972,358
Total Investment (his own money) $31,681
Ending Balance $972,358

Total Investment (his own money)

$0.032 million

Ending Balance

$0.972 million

C

If Sammy has contributed only for 23 years from 38 to 60 years
N A B=A*0.1 C=0.5*B D=B+C E=D*((1.125^(30-N)
Year Expected Salary Expected Saving Expected Match Total Saving FV of Total Saving
9 $46,088 $4,609 $2,304 $6,913 82013.44844
10 $47,701 $4,770 $2,385 $7,155 75452.37256
11 $49,371 $4,937 $2,469 $7,406 69416.18276
12 $51,099 $5,110 $2,555 $7,665 63862.88814
13 $52,887 $5,289 $2,644 $7,933 58753.85708
14 $54,738 $5,474 $2,737 $8,211 54053.54852
15 $56,654 $5,665 $2,833 $8,498 49729.26464
16 $58,637 $5,864 $2,932 $8,796 45750.92347
17 $60,690 $6,069 $3,034 $9,103 42090.84959
18 $62,814 $6,281 $3,141 $9,422 38723.58162
19 $65,012 $6,501 $3,251 $9,752 35625.69509
20 $67,288 $6,729 $3,364 $10,093 32775.63948
21 $69,643 $6,964 $3,482 $10,446 30153.58833
22 $72,080 $7,208 $3,604 $10,812 27741.30126
23 $74,603 $7,460 $3,730 $11,190 25521.99716
24 $77,214 $7,721 $3,861 $11,582 23480.23739
25 $79,916 $7,992 $3,996 $11,987 21601.8184
26 $82,714 $8,271 $4,136 $12,407 19873.67292
27 $85,609 $8,561 $4,280 $12,841 18283.77909
28 $88,605 $8,860 $4,430 $13,291 16821.07676
29 $91,706 $9,171 $4,585 $13,756 15475.39062
30 $94,916 $9,492 $4,746 $14,237 14237.35937
SUM $148,998 SUM $861,438
Total Investment (his own money) $148,998
Ending Balance $861,438

Total Investment (his own money)

$0.149 million

Ending Balance

$0.861 million

IF YOU HAVE ANY CLARIFICATIONS PLEASE ASK IN THE COMMENT BOX,I'AM HERE TO HELPS YOU.PLEASE GIVE ME THUMBS UP RATING

**********THANK YOU***********